Chickengirl1116:

Given that PS is the perpendicular bisector of QR, PQ=3.5m+18, and PR=6m+3, identify PQ. PQ = 39 PQ = 6 PQ = 17 PQ = 13

Lexxi:

PQ = PR 3.5m + 18 = 6m + 3 3.5m - 6m + 18 = 6m - 6m + 3 -2.5m + 18 = 3 -2.5m + 18 - 18 = 3 - 18 -2.5m = - 15 -2.5m/-2.5 = -15/-2.5 m = 6 PQ = 3.5m+18 = 3.5(6) + 18 = 21 + 18 = 39 PQ = 39